Good Morning!
I am trying so solve some problems.
The selling price of x number of a certain stereo can be modeled by the function R(x)=150x. The total cost of making x stereos is C(x)= 90x-0.02x^2. What is the percent markup for 60 stereos?
Thanks for the help!!
Denise

Answer by ewatrrr(24785)


Hi
Question states:
selling price of 60 of a certain stereo = 150*60 = $9000
The total cost of making 60 stereos = 90*60-0.02(60)^2 = 5400 - 72 = $5328
Percent markup is ($9000-$5328)/$5328 * 100 = 68.9 %